# Dynamic Pricing with Adversarially-Censored Demands

### v1

@@ 1 Introduction
The problem of dynamic pricing, where the seller proposes and adjusts their prices over time, has been studied since the seminal work of Cournot, (1897). The crux to pricing is to balance the profit of sales per unit with the quantity of sales. Therefore, it is imperative for the seller to learn customers’ demand as a function of price (commonly known as the demand curve) on the fly. However, the demand can often be obfuscated by the observed quantity of sales, especially when censored by inventory stockouts. Such instances severely impede the seller from learning the underlying demand distributions, thereby hindering our pursuit of the optimal price.
Existing literature has devoted considerable effort to the intersection of pricing and inventory decisions. Such works often consider scenarios with indirectly observable lost demands (Keskin et al.,, 2022), recoverable leftover demands Chen et al., (2019), or controllable inventory level (Chen et al.,, 2024). However,these assumptions do not always align with the realities faced in various common business environments. To illustrate, we present two pertinent examples:
Example 1.1 (Performance Tickets).
Imagine that we manage a touring company that arranges a series of performances featuring a renowned artist across various cities. Each venue has a different seating capacity, which substantially affects how we set ticket prices. If the price is too high, it might deter attendance, leading to lower revenue. On the other hand, setting it too low could mean that tickets sell out quickly, leaving many potential attendees unable to purchase them. We do not know exactly how many people attempt to buy tickets and fail. Moreover, because the performances are unique, there is no assurance that those who miss out on one show will choose or be able to attend another. This variability in venue size across different locations requires us to continually adapt our pricing strategy, ensuring that we maximize both attendance and revenue while accommodating unpredictable changes in seat availability.
Example 1.2 (Fruit Retails).
Sweetsop (Annona squamosa, or so-called "sugar apple") is a particularly-perishable tropical fruit, typically lasting only 2 to 4 days (Crane et al.,, 2005). Suppose we manage a local fruit shop and have partnered with a nearby farm for the supply of sweetsops during the harvest season. Due to the perishability nature, we receive sweetsops as soon as they are ripen and picked from the farm every day. This irregular supply means that some days we might receive a large quantity while getting very few on other days. We must quickly sell these fruits before they spoil, yet managing the price becomes challenging. If we exhaust our inventory ahead of time, customers will turn to other fruit shops to make their purchases rather than waiting for our next restock.
Products in the two instances above have the following properties,
@@
    Inventory level is determined by natural factors, and is arbitrarily given for every individual time period.
@@
    Products are perishable and salable only within a single time period.
@@
Dynamic pricing with adversarial inventory constraint. For $t=1,2,...,T:$ 1. The seller (we) receives $\gamma_{t}$ identical products. 2. The seller proposes a price $p_{t}\geq 0$. 3. The customers generate an invisible potential demand $d_{t}\geq 0$, dependent on $p_{t}$. 4. The market reveals an inventory-censored demand $D_{t}=\min\{\gamma_{t},d_{t}\}$. 5. The seller gets a reward $r_{t}=p_{t}\cdot D_{t}$. 6. All unsold products perish before $t+1$.
@@ 1.1 Summary of Contributions
We consider the problem setting displayed above and assume the potential demand $d_{t}=a-bp_{t}+N_{t}$ is linear and noisy. Here $a,b\in\mathbb{R}^{+}$ are fixed unknown parameters and $N_{t}$ is an unknown i.i.d. [^1] Independently and identically distributed noise with zero mean. Under this premise, the key to deriving the optimal price is to accurately learn the expected reward function $r(p)$, which is equivalent to learning the linear parameters $[a,b]$ and the noise distribution. We are confronted by three principal challenges:
@@
[…]
@@
To the best of our knowledge, we are the first to study the online dynamic pricing problem under adversarial inventory levels. Our C20CB algorithm attains an optimal $\tilde{O}(\sqrt{T})$ regret guarantee with high probability. The methodologies we develop are not only pivotal to our design and analysis but also potential for broad application in a variety of online decision-making scenarios with censored feedback.
@@ 1.2 Paper Structure
The rest of this paper is organized as follows. We discuss and compare with related works in Section 2, and then describe the problem setting in Section 3. We propose our main algorithm C20CB in Section 4 and analyze its regret guarantee in Section 5. We further discuss the limitations and potential extensions of our methodologies in Section 6, followed by a brief conclusion in Section 7.
@@ 2 Related Works
There exists a large volume of literature related to the problem we study in this work. Here we discuss them in the following categories.
Data-driven dynamic pricing
@@
Dynamic pricing for identical products is a well-established research area, starting with Kleinberg and Leighton, (2003) and continuing through seminal works by Besbes and Zeevi, (2009); Broder and Rusmevichientong, (2012); Wang et al., (2014, 2021). The standard approach involves learning a demand curve from price-sensitive demand arriving in real-time, aiming to approximate the optimal price. Kleinberg and Leighton, (2003) provided algorithms with regret bounds of $O(T^{\frac{2}{3}})$ and $O(\sqrt{T})$ for arbitrary and infinitely smooth demand curves, respectively. Wang et al., (2021) refined this further, offering an $O(T^{\frac{k+1}{2k+1}})$ regret for $k$-times continuously differentiable demand curves. This line of inquiry is also intricately linked to the multi-armed bandit problems (Lai and Robbins,, 1985; Auer et al.,, 2002) and continuum-armed bandits (Kleinberg,, 2004), where each action taken reveals a reward without insight into the foregone rewards of other actions.
Contextual pricing: Linear valuation and binary-censored demand
A surge of research has delved into feature-based dynamic pricing (Cohen et al.,, 2020) or pricing with contexts/covariates (Amin et al.,, 2014; Miao et al.,, 2019; Liu et al.,, 2021). These works considered situations where each pricing period was preceded by a context, influencing both the demand curve and noise distribution. Specifically, Cohen et al., (2020); Javanmard and Nazerzadeh, (2019); Xu and Wang, (2021) explored a linear valuation framework with known distribution noise, leading to binary customer demand outcomes based on price comparisons to their valuations. Expanding on this, Golrezaei et al., (2019); Fan et al., (2024); Luo et al., (2021); Xu and Wang, (2022) examined similar models but with unknown noise distributions. In another vein, Ban and Keskin, (2021); Wang et al., (2025); Xu and Wang, (2024) investigated personalized pricing where demand was modeled as a generalized linear function sensitive to contextual price elasticity. Many of these works on valuation-based contextual pricing also assume a censored demand: The seller only observes a binary feedback determined by a comparison of price with valuation, instead of observing the valuation directly. However, it was important to differentiate between the linear (potential) demand model we assumed and their linear valuation models, and there exists no inclusive relationship from each other.
Network Revenue Management (NRM)
NRM (Talluri and Van Ryzin,, 2006) studies the pricing and allocation problem on shared resources in a network. In the problem settings of Besbes and Zeevi, (2012) and the following works Simchi-Levi et al., (2019); Miao and Wang, (2024), if we take a marginal observation from each different product, the occupation of resources by other products may result in adversarial supply. In this work, we also consider adversarial inventories. However, our paper focuses on resolving the censoring effect on demand functions and reducing regret through online learning methodologies. Perakis and Roels, (2010) also adopts regret as a metric and considers censoring effect on the demand data as we do. However, their definition of minimax and maximin regret are different from ours. Also, instead of providing theoretical solutions, the censoring effect is only considered in empirical validation to test the robustness of the control system. Other representative works on NRM including Gallego and Van Ryzin, (1994) that proposes a classic model for dynamic pricing with stochastic demand, highlighting structural monotonicity and showing that simple policies can be asymptotically optimal; Talluri and Van Ryzin, (1998), which shows bid-price controls are near-optimal in large capacities but are not strictly optimal; and Meissner and Strauss, (2012) that addresses overbooking with product-specific no-shows via a randomized linear program.
Pricing with inventory concerns
@@
Dynamic pricing problems began to incorporate inventory constraints with the work of Besbes and Zeevi, (2009), which assumed a fixed initial stock available at the start of the selling period. They introduced near-optimal algorithms for both parametric and non-parametric demand distributions, operating under the assumption that the inventory was non-replenishable and non-perishable. Wang et al., (2014) adopted a comparable framework but allowed for customer arrivals to follow a Poisson process. In these earlier works, the actual demand is fully disclosed until the inventory is depleted. Subsequent research allowed inventory replenishment, with the seller’s decisions encompassing both pricing and restocking at each time interval. Chen et al., (2019) proposed a demand model subject to additive/multiplicative noise and developed a policy that achieved $O(\sqrt{T})$ regret. More recent studies, such as those by Chen et al., (2020); Keskin et al., (2022); Xu et al., (2025) explored the dynamic pricing of perishable goods where unsold inventory would expire. However, the uncensored demand is observable as assumed in both works. Specifically, Chen et al., (2020) allowed recouping backlogged demand, albeit at a cost, and introduced an algorithm with optimal regret. Keskin et al., (2022) focused on cases where both fulfilled demands and lost sales were observable.
@@
Chen et al., (2021) and their subsequent work, Chen et al., (2024), are the closest works to ours as they adopt similar problem settings: In their works, the demand is censored by the inventory level and any leftover inventory or lost sales disappear at the end of each period. With the assumption of concave reward functions and the restriction of at most $m$ price changes, Chen et al., (2021) proposed MLE-based algorithms that attain a regret of $\tilde{O}(T^{\frac{1}{m+1}})$ in the well-separated case and $\tilde{O}(T^{\frac{1}{2}+\epsilon})$ for some $\epsilon=o(1)$ as $T\rightarrow\infty$ in the general case. Under similar assumptions (except infinite-order smoothness), Chen et al., (2024) developed a reward-difference estimator, with which they not only enhanced the prior result for concave reward functions to $\tilde{O}(\sqrt{T})$ but also obtained a general $\tilde{O}(T^{2/3})$ regret for non-concave reward functions. Our problem model mirrors their difficulty in lacking access to both the uncensored demand and its gradient. However, they allowed the sellers to determine inventory levels with sufficient flexibility, hence better balancing the information revealed by the censored demand and the reward from (price, inventory) decisions. On the other hand, we assume the inventory level at each time period is provided adversarially by nature, which could impede us from learning the optimal price in the worst-case scenarios. Furthermore, due to the non-stationarity of inventory levels in our setting, the optimal price $p_{t}^{*}$ deviates over time. Given this, the searching-based methods adopted in Chen et al., (2024) is no longer applicable to our problem.
@@ 3 Problem Setup
In this work, we study the following online non-contextual dynamic pricing problem. At each time step $t=1,2,\ldots,T$, the seller (we) first receives an inventory quantity $\gamma_{t}$ determined by adversarial nature. Then we propose a price $p_{t}$ for these items and receive an inventory-censored demand $D_{t}:=\min\{d_{t},\gamma_{t}\}$. Here $d_{t}=a-b\cdot p_{t}+N_{t}$ is a potential linear demand, where $a,b$ are unknown parameters and $N_{t}$ is the demand noise.
In the following subsections, we will adopt definitions and assumptions to further clarify the scope of our methodologies.
@@ 3.1 Definitions
Here we define some key quantities that are involved in the algorithm design and analysis. Firstly, there are two different types of demand functions.
Definition 3.1 (Demand functions).
Denote $d_{t}(p):=a-bp+N_{t}$ as the potential demand function, and $d(p):=a-bp$ as the expected potential demand function. Denote $D_{t}(p):=\min\{\gamma_{t},d_{t}(p)\}$ as the censored demand function.
Then we define distributional functions of the noise $N_{t}$.
Definition 3.2 (Distributional functions).
For $N_{t}$ as the demand noise, denote $F(x)$ as its cumulative distribution function (CDF), $x\in\mathbb{R}$. Also, denote
[…]
as the integrated CDF.
We will make more assumptions on the noise distribution later. Notice that we do not assume the existence of PDF for $N_{t}$. However, if there exists its PDF in specific cases, we will adopt $f(x)$ as a notation.
Finally, we may define the revenue function and the regret.
Definition 3.3 (Revenue function).
@@ 3.2 Assumptions
We make reasonable and mild assumptions as follows. Firstly, we assume boundaries for parameters and price.
Assumption 3.5 (Boundedness).
There exist known finite constants $a_{\max},b_{\min},b_{\max},\gamma_{\min},c>0$ such that $0<a\leq a_{\max}$, $0<b_{\min}\leq b\leq b_{\max}$, $\gamma_{t}\geq\gamma_{\min}$, $N_{t}\in[-c,c]$. Also, we restrict the proposed price $p_{t}$ at any $t=1,2,\ldots,T$ satisfies $0\leq p_{t}\leq p_{\max}$ with a known finite constant $p_{\max}>0$.
Secondly, we make assumptions on the noise distribution.
Assumption 3.6 (Noise Distribution).
Each $N_{t}$ is drawn from an unknown independent and identical distribution (i.i.d.) satisfying $\mathbb{E}[N_{t}]=0$. The CDF $F(x)$ is $L_{F}$-Lipschitz. Also, according to Assumption 3.5, we have $F(-c)=0,F(c)=1$.
@@
Assumption 3.7 (Inequalities of Parameters).
@@
The parameters and constants satisfy the following inequalities:
  1. 1.
    $a-c>\gamma_{t},\forall t\in[T]$. Demands at $p_{t}=0$ must be censored.
  2. 2.
    $\gamma_{t}>2c,\forall t\in[T]$. Inventory level exceeds noise support.
  3. 3.
    $a-bp_{\max}-c>0$ and therefore $p_{\max}\leq\frac{a}{b}$. Demands must be positive.
  4. 4.
    $\gamma_{\min}>a_{\max}-b_{\min}p_{\max}+c$. Demands at $p_{t}=p_{\max}$ must be uncensored. Meanwhile, we denote $\gamma_{0}:=a_{\max}-b_{\min}p_{\max}+c$ for further use.
  5. 5.
@@
    Without loss of generality, let $p_{\max}\geq\frac{a}{2b}+1$. Optimal price must be included in $[0,p_{\max}]$.
Finally, we assume that the time horizon $T$ is sufficiently large, such that it will not confound with any constant or coefficient.
Assumption 3.8 (Large $T$).
For any given polynomial of parameters $poly(a_{\max},b_{\max},1/b_{\min},c,1/\gamma_{\min},p_{\max})$, we have $T>poly$.
4 Algorithm

### v2

@@ 1 Introduction
The problem of dynamic pricing, where the seller proposes and adjusts their prices over time, has been studied since the seminal work of [24]. The crux to pricing is to balance the profit of sales per unit with the quantity of sales. Therefore, it is imperative for the seller to learn customers’ demand as a function of price (commonly known as the demand curve) on the fly. However, the demand can often be obfuscated by the observed quantity of sales, especially when censored by inventory stockouts. Such instances severely impede the seller from learning the underlying demand distributions, thereby hindering our pursuit of the optimal price.
Existing literature has devoted considerable effort to the intersection of pricing and inventory decisions. Such works often consider scenarios with indirectly observable lost demands [33], recoverable leftover demands [12], or controllable inventory level [15]. However, these assumptions do not always align with the realities faced in various common business environments. To illustrate, we present two pertinent examples:
Example 1(Performance Tickets)
Imagine that we manage a touring company that arranges a series of performances featuring a renowned artist across various cities. Each venue has a different seating capacity, which substantially affects how we set ticket prices. If the price is too high, it may deter attendance, leading to lower revenue. On the other hand, setting it too low could mean that tickets sell out quickly, leaving many potential attendees unable to purchase them. We do not know exactly how many people attempt to buy tickets and fail. Moreover, because the performances are unique, there is no assurance that those who miss out on one show will choose or be able to attend another. This variability in venue size across different locations requires us to continually adapt our pricing strategy. With more adaptive prices, we can maximize both attendance and revenue while accommodating unpredictable changes in seat availability.
Example 2(Fruit Retails)
Sweetsop (Annona squamosa, or so-called “sugar apple”) is a particularly-perishable tropical fruit, typically lasting only 2 to 4 days [25]. Suppose we manage a local fruit shop and have partnered with a nearby farm for the supply of sweetsops during the harvest season. Due to their perishable nature, we receive sweetsops as soon as they are ripe and picked from the farm every day. This irregular supply means that some days we might receive a large quantity while getting very few on other days. We must quickly sell these fruits before they spoil, yet managing the price becomes challenging. If we exhaust our inventory ahead of time, customers will turn to other fruit shops for purchase instead of waiting for our next restock.
Products in the two instances above have the following properties:
@@
    Inventory levels are determined by natural factors, and are arbitrarily given for different individual time periods.
@@
    Products are perishable and only salable within a single time period.
1.1 Problem Overview
@@
Dynamic pricing with adversarial inventory constraint. For t=1,2,,T:t=1,2,...,T: 1. The seller (we) receives $\gamma_{t}$ identical products. 2. The seller proposes a price $p_{t}\geq 0$. 3. The customers generate an invisible potential demand $d_{t}\geq 0$, dependent on $p_{t}$. 4. The market reveals an inventory-censored demand $D_{t}=\min\{\gamma_{t},d_{t}\}$. 5. The seller gets a reward $r_{t}=p_{t}\cdot D_{t}$. 6. All unsold products perish before $t+1$.
The notion of “adversarial” inventory.
We characterize this problem as having adversarially-chosen inventory levels, though our setting differs from classical adversarial bandits or online convex optimization. Here, an adversary may pre-commit to an arbitrary inventory sequence $\{\gamma_{t}\}_{t=1}^{T}$, while the underlying demand noise remains stochastic—a hybrid model that combines adversarial contexts with stochastic outcomes, following the convention of “adversarial features” in online learning [21, 39]. Since inventory $\gamma_{t}$ is revealed at the beginning of round $t$, we can define per-round optimal prices $p_{t}^{*}$. This leads to a stronger regret benchmark: we compete against this sequence of optimal actions $\{p_{t}^{*}\}_{t=1}^{T}$ rather than the best fixed decision in hindsight typical of standard adversarial settings. This distinction is crucial as the inventory levels create a non-stationary optimization landscape where optimal prices vary dramatically across periods.
@@ 1.2 Summary of Contributions
We consider the problem setting shown above and assume that the potential demand $d_{t}=a-bp_{t}+N_{t}$ is linear and noisy. Here $a,b\in\mathbb{R}^{+}$ are fixed unknown parameters and $N_{t}$ is an unknown and i.i.d. (independently and identically distributed) noise with zero mean. Under this premise, the key to obtaining the optimal price is to accurately learn the expected reward function $r(p)$, which is equivalent to learning the linear parameters $[a,b]$ and the noise distribution. We are faced with three principal challenges:
@@
In this paper, we introduce an algorithm that employs innovative techniques to resolve the aforementioned challenges. Firstly, we devise a pure-exploration phase that bypasses the censoring effect and obtains an unbiased estimator of $\frac{1}{b}$ (which leads to $\hat{b}$ and $\hat{a}$ as a consequence). Secondly, we maintain estimates of the noise CDF $F(x)$ and $\int F(x)dx$ over a series of discrete $x$’s, as well as the confidence bounds of each estimate. Thirdly, we design an optimistic strategy, C20CB as “Closest-To-Zero Confidence Bound”, that proposes the price $p_{t}$ whose reward derivative $r^{\prime}_{t}(p_{t})$ is probably 0 or closest to 0 among a set of discretized prices. As we keep updating the estimates of $r^{\prime}_{t}(\cdot)$ with shrinking error bar, we asymptotically approach the optimal price $p^{*}_{t}$ since $r^{\prime}_{t}(p_{t}^{*})=0$ for any $t=1,2,\ldots,T$.
@@
To the best of our knowledge, we are the first to study the online dynamic pricing problem under adversarial inventory levels. Our C20CB algorithm attains an optimal $\tilde{O}(\sqrt{T})$ regret guarantee with high probability. The methodologies we develop are crucial to our algorithmic design, and are potentially advancing a variety of online decision-making scenarios with censored feedback.
@@ 1.3 Paper Structure
The rest of this paper is organized as follows. We discuss related works in Section˜2, and then describe the problem setting in Section˜3. We propose our main algorithm C20CB in Section˜4 and analyze its regret guarantee in Section˜5. We further discuss potential extensions in Section˜6, followed by a brief conclusion in Section˜7.
@@ 2 Related Works
Here we discuss the closest related works on dynamic pricing, inventory constraints, and network revenue management. For a broader introduction of related literature, please refer to Appendix˜0.A.
Data-driven dynamic pricing.
@@
Dynamic pricing for identical products is a well-established research area, starting with [36] and continuing through seminal works by [7, 10, 57, 55]. The standard approach involves learning a demand curve from price-sensitive demand arriving in real-time, aiming to approximate the optimal price. [36] provided algorithms with regret bounds of $O(T^{\frac{2}{3}})$ and $O(\sqrt{T})$ for arbitrary and infinitely smooth demand curves, respectively. [55] refined this further, offering an $O(T^{\frac{k+1}{2k+1}})$ regret for $k$-times continuously differentiable demand curves. This line of inquiry is also intricately linked to the multi-armed bandit problems [37, 4] and continuum-armed bandits [35], where each action taken reveals a reward without insight into the foregone rewards of other actions.
Pricing with inventory concerns.
@@
Dynamic pricing problems begin to incorporate inventory constraints with [7], which assumed a fixed initial stock available at the start of the selling period. They introduced near-optimal algorithms for both parametric and non-parametric demand distributions, operating under the assumption that the inventory is non-replenishable and non-perishable. [57] adopted a comparable framework but allowed customers arrivals to follow a Poisson process. In these earlier works, the actual demand is fully disclosed until the inventory is depleted. Subsequent research allows inventory replenishment, with the seller’s decisions encompassing both pricing and restocking at each time interval. [12] proposed a demand model subject to additive / multiple noise and developed a policy that achieved $O(\sqrt{T})$ regret. More recent studies [14, 33] explored the pricing of perishable goods where the unsold inventory will expire. However, the uncensored demand is observable as assumed in both works. Specifically, [14] allowed recouping backlogged demand, albeit at a cost, and introduced an algorithm with optimal regret. [33] focused on the cases where both fulfilled demands and lost sales were observable.
@@
[13] and their subsequent work, [15], are the closest works to ours as they adopt similar problem settings: In their works, the demand is censored by the inventory level and any leftover inventory or lost sales disappear at the end of each period. With the assumption of concave reward functions and the restriction of at most $m$ price changes, [13] proposed MLE-based algorithms that attain a regret of $\tilde{O}(T^{\frac{1}{m+1}})$ in the well-separated case and $\tilde{O}(T^{\frac{1}{2}+\epsilon})$ for some $\epsilon=o(1)$ as $T\rightarrow\infty$ in the general case. Under similar assumptions (except infinite-order smoothness), [15] developed a reward-difference estimator, with which they not only enhanced the prior result for concave reward functions to $\tilde{O}(\sqrt{T})$ but also obtained a general $\tilde{O}(T^{2/3})$ regret for non-concave reward functions. Our problem model mirrors their difficulty, as we also lack access to both the uncensored demand and its gradient. However, they allowed the sellers to determine inventory levels with sufficient flexibility, hence better balancing the information revealed by the censored demand and the reward from (price, inventory) decisions. On the other hand, we assume that the inventory level at each time period is provided adversarially by nature, which could impede us from learning the optimal price in the worst-case scenarios. Furthermore, due to the non-stationarity of inventory levels in our setting, the optimal price $p_{t}^{*}$ deviates over time. Given this, the search-based methods adopted in [15] are no longer applicable to our problem.
Network Revenue Management (NRM).
NRM [50] studies pricing and allocation of shared resources in a network. In the settings of [8] and subsequent works [47, 44], marginal observations for each product can induce adversarial supply due to cross-product resource occupation. We also consider adversarial inventories, but focus on addressing demand censoring and reducing regret via online learning. [45] likewise used regret and considered censoring in demand data, but their minimax/maximin regret definitions differ from ours, and censoring is mainly used in empirical validation rather than for theoretical solutions. Representative NRM works include [28] on dynamic pricing with stochastic demand, establishing structural monotonicity and asymptotic optimality of simple policies; [49] showing bid-price controls are near-optimal in large capacities but not strictly optimal; and [42] addressing overbooking with product-specific no-shows via a randomized LP.
@@ 3 Problem Setup
We have defined the problem setting in Section˜1.1. To further clarify the scope of our methodology, we make the following very first assumption before introducing further concepts.
Assumption 1 (Linear Demand)
Assume the potential demand $d_{t}=d_{t}(p):=a-bp+N_{t}$ is linear and noisy. Denote d(p):=abpd(p):=a-bp as the expected potential demand function. Denote $D_{t}(p):=\min\{\gamma_{t},d_{t}(p)\}$ as the censored demand function.
A linear demand model has been widely used in pricing literature, including [40, 46, 10, 23]. We highlight two primary reasons: (1) According to [9], linear demand allows the prices to converge to the true optimum even under mismatching scenarios. (2) As shown in [34], the same analysis can be applied to Generalized Linear Model (GLM), which has enhanced capability of capturing the real-world demands (see e.g. [11, 54]).
@@ 3.1 Definitions
Here we define some key quantities that are involved in the algorithm design and analysis. Firstly, we define distributional functions of the noise $N_{t}$.
Definition 1(Distributional Functions)
For $N_{t}$ as the demand noise, denote $F(x)$ as its cumulative distribution function (CDF), $x\in\mathbb{R}$. Also, denote the following $G(x)$ as the integrated CDF:
$G(x):=\int_{-\infty}^{x}F(\omega)d\omega,x\in\mathbb{R}$ (1)
We will make more assumptions on the noise distribution later. Notice that we do not assume the existence of PDF for $N_{t}$. However, if there exists its PDF in specific cases, we will adopt $f(x)$ as a notation. Then, we define the revenue function and the regret.
Definition 2(Revenue Function)
@@
The definition of regret inherits from the tradition of online learning, capturing the performance difference between the algorithm-in-use and the best benchmark that an omniscient oracle can achieve (which knows everything except the realization of noise).
@@ 3.2 Assumptions
Firstly, we assume boundaries for parameters and price.
Assumption 2 (Boundedness)
There exist known finite constants $a_{\max},b_{\min},b_{\max},\gamma_{\min},$ and $c>0$ such that $0<a\leq a_{\max}$, $0<b_{\min}\leq b\leq b_{\max}$, $\gamma_{t}\geq\gamma_{\min}$, $N_{t}\in[-c,c]$. Also, we restrict the proposed price $p_{t}$ at any $t=1,2,\ldots,T$ satisfies $0\leq p_{t}\leq p_{\max}$ with a known finite constant $p_{\max}>0$.
The assumption of boundedness on $a,b$ and price is natural as it defines the scope of instances. We justify the assumption of noise with bounded support [c,c][-c,c] from two aspects: (1) The upper bound exists without loss of generality due to the existence of inventory-censoring effect. (2) The lower bound exists as we avoid negative demand. Secondly, we make assumptions on the noise distribution.
Assumption 3 (Noise Distribution)
Each $N_{t}$ is drawn from an unknown independent and identical distribution (i.i.d.) satisfying $\mathbb{E}[N_{t}]=0$. The CDF $F(x)$ is $L_{F}$-Lipschitz. Also, according to Assumption˜2, we have F(c)=0,F(c)=1F(-c)=0,F(c)=1.
@@
Assumption 4 (Inequalities of Parameters)
@@
The parameters and constants involved in the problem setting satisfy the following conditions:
Assumption Explanation
(1) $a-c>\gamma_{t},\forall t\in[T]$
The demand at $p=0$ is completely censored by any inventory level,
since customers will rush to buy until completely out-of-stock.
(2) $\gamma_{t}>2c,\forall t\in[T]$
Inventory level should exceed the width of the noise support.
Otherwise we can reshape the noise by capping $N_{t}$ at $\gamma_{t}-a+bp_{\max}$.
(3) $a-bp_{\max}-c>0$ Demands must be positive.
(4) $\gamma_{\min}>a_{\max}-b_{\min}p_{\max}+c$
Demands at $p_{t}=p_{\max}$ must be uncensored.
We denote $\gamma_{0}:=a_{\max}-b_{\min}p_{\max}+c$ for further use.
@@
(5) $p_{\max}\geq\frac{a}{2b}$ Optimal price must be included in $0,p_{\max}]$ without loss of generality.
Each assumption in Assumption˜4 is justified by an explanation followed. We make Item (3) to avoid simultaneously upper-and lower-censoring effect, which is beyond the scope of this work and considered as an extension. It is worth mentioning that for those boundedness parameters (e.g. $a_{\max},b_{\max},c,\gamma_{\min}$), we have to know their exact values, while we only require the existence of $a,b$ without knowing them. Finally, for the benefit of regret analysis, we assume that the time horizon $T$ is sufficiently large, such that its polynomial will not confound any constant or coefficient.
Assumption 5 (Large $T$)
For any constant $r=O(1)$, time horizon $T$ is larger than any polynomial of parameters, i.e. $T>\Omega\left((\frac{a_{\max}b_{\max}cp_{\max}}{b_{\min}\gamma_{\min}})^{r}\right)$.
4 Algorithm Design